OpenStudy (anonymous):

Using the formula,, find the speed of a car in miles per hour if it took 150 feet to stop. Round your answer to a whole number. Answer 40 mph 60 mph 75 mph

OpenStudy (anonymous):

@JennWoww

OpenStudy (anonymous):

yes @Chooch146 ?

OpenStudy (anonymous):

s = sqrt (24d) s^2 = 24 * 150 s^2 = 3600 s = 60

OpenStudy (anonymous):

\[\sqrt{24d}\]

OpenStudy (anonymous):

can i get medal??

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

lol your welcome:]